OpenStudy (anonymous):

cos(a+b)/(cos(a)cos(b)=1-tan(a) tan(b) verify the identity step by step

OpenStudy (kirbykirby):

Try replacing these in your equation: \(\cos(a+b)=\cos (a)\cos (b)-\sin (a)\sin (b )\) \(\tan( a) =\dfrac{\sin(a)}{\cos(a)}\), and do the same for \(\tan(b)\) You should be able to seequickly how both sides are equal when you do this

OpenStudy (anonymous):

sounds good. Thank you!

OpenStudy (kirbykirby):

=] your welcome